OpenStudy (anonymous):

How many atoms are in 20.9 grams of Au?

OpenStudy (anonymous):

For this question you have to convert the grams of Au into moles by dividing it by the molar mass of Au, then multiply it times Avogadro's number in order to obtain the atoms.\[20.9g \times \frac{ 1 mol }{ 196.96655g } \times \frac{ 6.0221413e+23 atoms }{ 1 mol }\]\[=6.39 \times 10^{23} atoms of Au\]